5. An AM transmitter has a carrier power of 30 W. The percentage of modulation is 85 Calculate the following: (a)
Answer #1

The calculation for the problem is as follows:

1. Given that: Pc=30 W

2. Modulation index m= 0.85

(a) Total power in any AM signal is given by:

Pt =Pc (1+m2/2)

Substituting values:

Pt = 30(1+0.852/2)

    = 30(1+0.36125)

   =30*1.36125

    =40.8375W

Hence total power is: 40.8375W

(b)Power in side bands

Now total power is: 40.8375W

So total power in two side bands is 40.8375-30=10.8375W

So the power in each side band is 10.8375/2=5.41875W